OpenStudy (anonymous):

A function is defined as f(x) = 2x + 3 + 64/x^2. Find the values of x for which the function is increasing. So, I differentiated it and I get f'(x) as 2 - 128/x^3. For the function to be increasing the gradient must be positive. so 2 - 128/x^3 > 0. Solving this I can get x > 4. However, it's obvious that it also works when x < 0. (Taking -1, -2 - (-128/1) = -2 + 128 = 126. Which is greater than 0. What am I doing that makes me lose an answer? Thanks.

OpenStudy (anonymous):

2 + 128 = 130, which is greater than 0. I put an extra minus in there but the problem's the same.

OpenStudy (anonymous):

is it all over x^2?

OpenStudy (anonymous):

\[f(x)=2x+3+\frac{64}{x^2}\]?

OpenStudy (anonymous):

Just the 64 over the x^2.

OpenStudy (anonymous):

if so you get \[f'(x)=2-\frac{128}{x^3}=\frac{2(x^3-64)}{x^3}\] yes?

OpenStudy (anonymous):

positive if \[x<0\] or \[x>4\]

OpenStudy (anonymous):

oh i see i should have read your question more carefully. yes you cannot ignore the denominator. it counts too

OpenStudy (anonymous):

you get \[\frac{2(x-4)(x^2+4x+8)}{x^3}\]

OpenStudy (anonymous):

so critical points at 4 and 0 it is negative between the critical points and positive outside them

OpenStudy (anonymous):

you cannot just look at the factor x - 4 you also have to consider x^3 as a factor, which is positive if x is and negative if x is negative

OpenStudy (anonymous):

this is clear yes?

OpenStudy (anonymous):

@alchemista i am working with the derivative

OpenStudy (anonymous):

Where did you get 2(x−4)(x2+4x+8)\x3 from?

OpenStudy (anonymous):

your derivative is \[f'(x)=2-\frac{128}{x^3}=\frac{2x^3-128}{x^3}=\frac{2(x^3-64)}{x^3}\] yes?

OpenStudy (anonymous):

in order to see where this is positive and negative we need to write over one denominator

OpenStudy (anonymous):

Yes. I did 2- 128/x^3 > 0 2x^3 - 128 > 0 x^3 > 64

OpenStudy (anonymous):

then i just factored \[x^3-64\] as difference of two cubes

OpenStudy (anonymous):

oh no \[2-\frac{128}{x^3}>0\] not the same as \[2x^3-128>0\]

OpenStudy (anonymous):

because \[2-\frac{128}{x^3}\neq 2x^3-128\]

OpenStudy (anonymous):

you have to make the fraction first and write \[\frac{2x^3-128}{x^3}>0\] and consider both numerator and denominator

OpenStudy (anonymous):

clear or no?

OpenStudy (anonymous):

Ok, so where do you go from there?